OpenStudy (anonymous):

Sol is building a planting bed in the shape of a trapezoid. The bed will be 9 in deep. The bases of the trapezoid are 2ft and 8 ft and the sides are both 5 ft long. How many cubic feet of topsoil is needed to fill the bed?

OpenStudy (amistre64):

area of a trap is (b+b)h/2 so times that by 9 after calculating for the height

OpenStudy (amistre64):

and if course converting to inches.... or viceaversa

OpenStudy (anonymous):

ha right on... I am trying to work it out now... and see ...

OpenStudy (anonymous):

base and base is 2 and 8 ?

OpenStudy (anonymous):

and the height is the 9?

OpenStudy (amistre64):

one thing to remember is that a 3-4-5 triangle is a rt trianlge; and they love to use it....

OpenStudy (amistre64):

5^2 = 3^2 + h^2 25 - 9 = h^2 16 = h^2 ; h=4

OpenStudy (anonymous):

good grief

OpenStudy (amistre64):

(2+8)4 ------ = area = 20 2

OpenStudy (amistre64):

9 inches aint a foot; so convert it; 9/12 = 3/4 20(3/4) = 5(3) = 15 cubic feet

OpenStudy (anonymous):

so 15 cubic feet of top soil is needed... ok.. let me finish my problem and check it to yours.... thank you again!

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

Alright... I get the same... but it would have been wrong... I didnt know to convert. Good beans~

OpenStudy (amistre64):

you get a ruler .... :) 1 ft = 12 in 9in(1ft) ------ = (9/12) ft = (3/4)ft 12 in

OpenStudy (anonymous):

I like how you show the work and where it all comes from... Very great on the clarification!!! :)

OpenStudy (anonymous):

yeah it totally made sense